# SpyChain: Multi-Vector Supply Chain Attacks on Small Satellite Systems

### v1

@@ 1 Introduction
  • New Satellite Malware Technique: We mapped all tactics, techniques, and procedures from a successful SpyChain deployment to the respective standardized strategies listed in the Space Attack Research and Tactic Analysis (SPARTA) matrix [63]. Our investigation found a novel execution technique, using multi-component coordination to achieve a sophisticated stealth attack. After notifying the maintainers, this technique is slated for inclusion in the upcoming iteration of the SPARTA matrix.
@@ 5.5 Identifying Novel Strategies
Developing SpyChain involved utilizing multiple tactics, techniques, and procedures (TTP). After defining our attack pipeline, we mapped each TTP used over each stage of the attack with the standardized SPARTA matrix [63]. Table 3 presents the complete list of TTPs that we utilized, showing that our attack methodology aligns with realistic attack scenarios. Although the list is comprehensive and accurately represents our design decisions, it overlooks a vital technique: the colluding behavior of compromised components. This novel technique is key to our findings because it allows SpyChain to bypass testing and detection. By distributing malicious logic, SpyChain avoids centralized signatures, increases persistence, and achieves a higher level of stealth. We successfully utilized this technique in Scenario 3, Scenario 4, and Scenario 5, demonstrating its effectiveness in sustaining covert activity during mission operations. We informed the SPARTA team of this colluding-component technique, and they confirmed it will be incorporated into the next iteration of the matrix.
@@ I. Access IA-0010 Unauthorized Access During Safe-Mode
Defense Evasion TBA Component Collusion for Evasion
Table 3: SPARTA techniques involved in SpyChain. Each technique corresponds to a specific tactic, technique, or procedure (TTP) used in the adversarial model. Recon. stands for Reconnaissance. I. Access stands for Initial Access. The last row is our proposed TTP and will be available in the next SPARTA Matrix release.
@@ 8 Conclusion
Open Science
Artifacts Provided. To support evaluation of our work, we make the following artifacts available:
  • Pseudocode and diagrams illustrating the structure of our malicious and defensive cFS components, including communication flow, trigger conditions, and mitigation logic.
  • Screenshots and representative logs from the NOS3 simulation and COSMOS ground station, showing telemetry traces and attack/defense outcomes that correspond to figures and tables in the paper.
  • Documentation and methodology sufficient for reproducing the scenarios in principle using the public NOS3 release, which is already open source and widely available.
Access During Review. At submission time, pseudocode, screenshots, and example logs will be hosted in an anonymized repository with links provided here: https://anonymous.4open.science/r/SpyChain-Artifact-BCE1. These materials are scrubbed of identifying metadata and are sufficient for reviewers to verify the plausibility of our design, methodology, and reported results.
Limitations and Responsible Disclosure. We deliberately do not release the complete source code of the malicious components. Because many operational spacecraft use cFS or closely related software, providing complete, flight-ready malware would pose a direct security risk. Instead, we provide high-level pseudocode and partial interfaces that illustrate functionality without enabling immediate weaponization. Defensive modules and mitigation strategies are fully described.
Similarly, we do not redistribute NOS3 itself, as it is already available as an open-source NASA project; our contribution lies in how it can be configured and extended, not in modifying or repackaging the simulator.
Post-Acceptance Plan. If the paper is accepted, we are open to sharing full component code and experiment scripts with trusted collaborators under controlled conditions, with the aim of improving defenses for flight software. We believe this balances reproducibility with our responsibility to prevent the distribution of weaponizable code.

### v2

@@ 1 Introduction
  • New Satellite Malware Technique: We mapped all tactics, techniques, and procedures from a successful SpyChain deployment to the respective standardized strategies listed in the Space Attack Research and Tactic Analysis (SPARTA) matrix [28]. Our investigation found a novel execution technique, using multi-component coordination to achieve a sophisticated stealth attack. After notifying the maintainers, this technique has been added to the SPARTA matrix.
@@ 5.5 Identifying Novel Strategies
Developing SpyChain involved utilizing multiple tactics, techniques, and procedures (TTP). After defining our attack pipeline, we mapped each TTP used over each stage of the attack with the standardized SPARTA matrix [28]. Table 3 presents the complete list of TTPs that we utilized, showing that our attack methodology aligns with realistic attack scenarios. Although the list is comprehensive and accurately represents our design decisions, it overlooks a vital technique: the colluding behavior of compromised components. This novel technique is key to our findings because it allows SpyChain to bypass testing and detection. By distributing malicious logic, SpyChain avoids centralized signatures, increases persistence, and achieves a higher level of stealth. We successfully utilized this technique in Scenario 3, Scenario 4, and Scenario 5, demonstrating its effectiveness in sustaining covert activity during mission operations. We informed the SPARTA team of this colluding-component technique, and they have incorporated it into the matrix as Component Collusion under Defense Evasion with ID DE-0012.
@@ I. Access IA-0010 Unauthorized Access During Safe-Mode
Defense Evasion DE-0012 Component Collusion
Table 3: SPARTA techniques involved in SpyChain. Each technique corresponds to a specific tactic, technique, or procedure (TTP) used in the adversarial model. Recon. stands for Reconnaissance. I. Access stands for Initial Access. The last row is our new TTP, Component Collusion.